OpenStudy (anonymous):

Is there a calculator that can factor for all the zeros in a function? Like online. I don't have a calculator.

OpenStudy (anonymous):

go to google

OpenStudy (anonymous):

I have, I'm not finding anything.

OpenStudy (anonymous):

try wolframalpha

OpenStudy (anonymous):

what exactly are you trying to find?

OpenStudy (anonymous):

Find the zeros of the function f(x) = x^2 +9x + 20 by factoring. I have a ton of these equations.

OpenStudy (anonymous):

Or functions I guess.

OpenStudy (anonymous):

oooh then factor!!

OpenStudy (anonymous):

\[x^2+9x+20=(x+4)(x+5)\]

OpenStudy (anonymous):

and so if \(x+4=0\iff x=-4\) and \(x+5=0\iff x=-5\)

OpenStudy (anonymous):

Where did the 4 and 5 come from?

OpenStudy (anonymous):

\(4\times 5=20\) and \(4+5=9\) so \[(x+4)(x+5)=x^2+4x+5x+4\times 5=x^2+9x+20\]

OpenStudy (anonymous):

oh wow. Math is weird. I get it now, thank you.

OpenStudy (anonymous):

if your job is to learn how to factor, i would not use a calculator it will not help at all

OpenStudy (anonymous):

yeah, math is weird but it is not that weird if you can multiply out you can usually factor also these problems will be cooked up for you to factor easily